OpenStudy (flvskidd):

Determine if triangle RST with coordinates R (3,4), S (5,5) and T (6,1) is a right triangle. Use evidence to support your claim. If it is not a right triangle, what changes can be made to make it a right triangle? Be specific.

OpenStudy (anonymous):

find slope by formula \[m=\frac{ y _{2}-y _{1} }{x _{2}-x _{1}}\] taking two points at a time and see if m1*m2=-1

OpenStudy (flvskidd):

So I have to figure out if all possibilities of those points when subtracted equal to -1?

OpenStudy (anonymous):

slope of RS \[m _{1}=\frac{ 5-4 }{ 5-3 }=\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

slope of RT \[m _{2}=\frac{ 1-4 }{ 6-3 }=\frac{ -3 }{ 3 }=-1\]

OpenStudy (anonymous):

slope of ST \[m _{3}=\frac{ 1-5 }{ 6-5 }=\frac{ -4 }{ 1 }=-4\]

OpenStudy (flvskidd):

Now I see what you mean. This was very helpful.

OpenStudy (flvskidd):

So I would say that RS and ST need to equal -1 for it to be a right triangle?

OpenStudy (flvskidd):

Am I right? I will give you best answer.

OpenStudy (anonymous):

product of any two should be -1 which is not the case so change co-ordinates of one point to make the product -1

OpenStudy (flvskidd):

Product of any two? Okay you saved me there. Thank you very much.

OpenStudy (flvskidd):

Can you help me with another one?

OpenStudy (anonymous):

 adjust the coordinates

OpenStudy (anonymous):

2. find the slopes of CB and FE

OpenStudy (flvskidd):

For line FE I got -4/-5 and for line CB I got -3/-4

OpenStudy (flvskidd):

How do I know if they are opposite and reciprocal?

OpenStudy (anonymous):

slope of FE \[=\frac{ -4-(-1) }{ 5-1 }=\frac{ -3 }{ 4 }\] slope of CB \[=\frac{ 5-1 }{ 4-1 }=\frac{ 4 }{ 3 }\]